OpenStudy (anonymous):

derive this function: ln((sqrtx^2+1)/(x(2x^3-1)))

OpenStudy (dumbcow):

looks like you should use quotient rule (f/g)' = (f'g -fg') / g^2

OpenStudy (anonymous):

\[\ln (\sqrt{x ^{2}-1}\div x(2x ^{3}-1)^{2})\]

OpenStudy (anonymous):

i tried and it gets really complicated and long... i wondered if there's a simpler method

OpenStudy (anonymous):

just break it into two natural logs

OpenStudy (anonymous):

oh duh. thank you

OpenStudy (dumbcow):

oh the whole thing is inside the log...yeah log(x/y) = logx -logy then use chain rule